# Gravitational Wave Polarization Detection with Pyramid Constellation of Gravitational Wave Observatory

### v2

@@ 0.3 Polarization Modes
$\mathbf{H}(t)=h_{+}(t)\boldsymbol{\epsilon}^{+}+h_{\times}(t)\boldsymbol{\epsilon}^{\times}+h_{U}(t)\boldsymbol{\epsilon}^{U}+h_{V}(t)\boldsymbol{\epsilon}^{V}+h_{B}(t)\boldsymbol{\epsilon}^{B}+h_{L}(t)\boldsymbol{\epsilon}^{L},$ (4)
where $\boldsymbol{\epsilon}^{A}$ ($A=+,\times,U,V,B,L$) are the polarization tensors and $h_{A}$ are the waveforms of the polarization modes.
Solar System Barycenter ($SSB$) frame, based on the ecliptic plane, is selected. In $SSB$ frame, the standard spherical coordinates $\left(\theta,\phi\right)$ and the associated spherical orthonormal basis vectors $\left(\bf{e}r,\bf{e}\theta,\bf{e}\phi\right)$ are confirmed. The position of the source in the sky is parametrized by the ecliptic latitude $\beta=\pi/2-\theta$ and the ecliptic longitude $\lambda=\phi$. The gravitational wave propagation vector $\hat{k}$ in spherical coordinates is expressed as
@@
$\boldsymbol{\epsilon}^{+}$ $=$ $\mathbf{e}^{+}\cos 2\psi-\mathbf{e}^{\times}\sin 2\psi,$
$\boldsymbol{\epsilon}^{\times}$ $=$ $\mathbf{e}^{+}\sin 2\psi+\mathbf{e}^{\times}\cos 2\psi,$
$\boldsymbol{\epsilon}^{U}$ $=$ $\mathbf{e}^{U}\cos\psi-\mathbf{e}^{V}\sin\psi,$
$\boldsymbol{\epsilon}^{V}$ $=$ $\mathbf{e}^{U}\sin\psi+\mathbf{e}^{V}\cos\psi,$
$\boldsymbol{\epsilon}^{B}$ $=$ $\mathbf{e}^{B},$
$\boldsymbol{\epsilon}^{L}$ $=$ $\mathbf{e}^{L},$ (7)
where $\psi$ is the polarization angle and the six basis tensors are [7]
@@
{addendum}
@@
We thank Yue-Liang Wu for the theory of GW polarization beyond GR and requirement of the space mission of GW detection. which are value useful to improve the manuscript. This work has been supported by the Fundamental Research Funds for the Central Universities. This work has been supported in part by the Strategic Priority Research Program of the Chinese Academy of Sciences under Grant No. XDA15020701 and XDA15020708. In this paper, The numerical computation of this work was completed at TAIJI Cluster of University of Chinese Academy of Sciences.

### v3

@@ II.4 Polarization Modes
$\mathbf{H}(t)=h_{+}(t)\bm{\epsilon}^{+}+h_{\times}(t)\bm{\epsilon}^{\times}+h_{U}(t)\bm{\epsilon}^{U}$
$+h_{V}(t)\bm{\epsilon}^{V}+h_{B}(t)\bm{\epsilon}^{B}+h_{L}(t)\bm{\epsilon}^{L},$ (4)
where $\bm{\epsilon}^{A}$ ($A=+,\times,U,V,B,L$) are the polarization tensors and $h_{A}$ are the waveforms of the polarization modes.
Solar System Barycenter ($SSB$) frame, based on the ecliptic plane, is selected. In $SSB$ frame, the standard spherical coordinates $\left(\theta,\phi\right)$ and the associated spherical orthonormal basis vectors $\left(\bf{e}r,\bf{e}\theta,\bf{e}\phi\right)$ are confirmed. The position of the source in the sky is parametrized by the ecliptic latitude $\beta=\pi/2-\theta$ and the ecliptic longitude $\lambda=\phi$. The GW propagation vector $\hat{k}$ in spherical coordinates is expressed as
@@
$\bm{\epsilon}^{+}$ $=$ $\mathbf{e}^{+}\cos 2\psi-\mathbf{e}^{\times}\sin 2\psi,$
$\bm{\epsilon}^{\times}$ $=$ $\mathbf{e}^{+}\sin 2\psi+\mathbf{e}^{\times}\cos 2\psi,$
$\bm{\epsilon}^{U}$ $=$ $\mathbf{e}^{U}\cos\psi-\mathbf{e}^{V}\sin\psi,$
$\bm{\epsilon}^{V}$ $=$ $\mathbf{e}^{U}\sin\psi+\mathbf{e}^{V}\cos\psi,$
$\bm{\epsilon}^{B}$ $=$ $\mathbf{e}^{B},$
$\bm{\epsilon}^{L}$ $=$ $\mathbf{e}^{L},$ (7)
where $\psi$ is the polarization angle and the six basis tensors are (Nishizawa et al., 2009)
@@
II.5 Detector Response
The laser link signal is emitted from the S/C $i$ at $t_{i}$ moment and received by the S/C $j$ at $t_{j}$ moment. Detector arm Lij is the laser link from S/C $i$ to S/C $j$, whose displacement is $\ell_{ij}(t)$ at t moment. The gravitational wave travels in the $\hat{k}$ direction and ${\mathbf{x}(t)}$ is the position on the laser link at time t. By the Eq. (9), the $\delta\ell_{ij}(t)$ is calculated in a numerical way in the time domainGuo et al. (2024).
[…]
where, $\hat{r}_{ij}(t)$ denotes the unit vector
$\hat{r}_{ij}(t)=\frac{{\bf x}_{j}(t_{j})-{\bf x}_{i}(t_{i})}{\ell_{ij}(t)}\,.$ (10)
and ${\bf h}(t-\hat{k}\cdot{\mathbf{x}(t)})$ is the GW tensor in the transverse-traceless gauge, whose components are $\bm{h}_{A}$ ($A=+,\times,U,V,B,L$) from the Eq. (II.4). The amplitude of GW, responded to detector arm Lij, is in the following equation,
$\mathbf{H}_{ij}(t)=\frac{\delta\ell_{ij,A}(t)\epsilon^{A}}{\ell_{ij}(t)},$ (11)
where $\bm{\epsilon}^{A}$ ($A=+,\times,U,V,B,L$) are the polarization tensors from the Eq. (II.4). The upper and lower indicator A of the two quantities: ${\delta\ell_{ij,A}}$ and ${\epsilon^{A}}$ are summed. In the Fig. 5, $h_{sum}$ for Lij is $\mathbf{H}_{ij}(t)$ and $h_{A}$ for Lij is a component value equal to ${\delta\ell_{ij,A}(t)\epsilon^{A}}/{\ell_{ij}(t)}$ ( the upper and lower indicator A is only used as a component and not summed).
[…]
III Results
Comparing to those triangular detectors, TEGO has more than four times combinations of optical paths of time delay interferometer(TDI) to suppress the frequency noise. In this paper, based on the unequal arm Michelson TDI configuration and the Sagnac TDI configuration for three S/Cs, the optical path differences of the second-generation TDI are calculated to check the capability of TDI for the apparently variable arm lengths. As is seen in Figure 2, Michelson TDI configuration is from the optical path combination of two arms and the optical path differences are below the limits of noise canceling, compared to the results in the paper(Wang and Ni, 2019). In the optical path differences of all the combination, the values relevant to the combination between S/C1, S/C2 and S/C3 are 1% less than the others, which is seen in the Figure 2. This result is due to the fact that, over a one-year orbital period, three arm lengths relevant to S/C4 are changed apparently, that is seen in the right of Figure 1.
[…]
Figure 5: Amplitude of six polarization modes of GW source: J0806 response to all the arms of TEGO over a one-year orbital period. As an example, $L_{12}$ means the arm between S/C1 and S/C2.
In summary, as the radial distance of S/C4 is longer than the other S/Cs, three arm lengths relevant to S/C4 are changed apparently over a one-year orbital period. The second-generation TDI are selected to eliminate the laser frequency noise, instead of the first generation TDI. Comparing to the optical path differences shown in Figure 2, 3 and 4, the unequal arm Michelson TDI configuration and the Sagnac TDI configuration for three S/Cs are equally effective for elimination of the laser frequency noise.
A verification white dwarf binary signal: J0806 as the identified source is selected, whose parameters are detailed in the paper(Kupfer et al., 2018). The polarization angle is set $\psi=0$, and initial orbital phase is set $\phi_{0}=0$. The sky location parameters $\phi[rad],\beta[rad]$ and $D_{L}[kpc]$ are 2.1021, -0.0821 and 5 respectively(Kupfer et al., 2018). As a monochromatic source, The GW waveform of J0806 is calculated using the quadrupole approximation (Landau and Lifshitz, 1962; Peters and Mathews, 1963). Based on this approximation, the GW signals are described as a combination of the two polarizations (+,×+$\times$+ , ×)(Korol et al., 2022) and the other four components of GW polarizations cannot be calculated with that approximation. In this paper, the GW waveforms relevant to the other four components of GW polarizations are set to be equal to ($\times$) component, artificially. That is an independent model. The details are found in the subsection of Method section: GW Signal Model II.3. The every components of GW polarization expression in the subsection of Method section: Polarization Modes II.4.
As seen in Figure 5, all the detector arms are simultaneously sensitive to the six polarization modes of GWs. As the angle between a detector arm and the orientation of a polarization mode changes with the detector’s orbital position, a polarization mode of GWs exhibits variable amplitude over a one-year orbital period. This implies that there is an optimal orbital position for the detector where the amplitude of a given polarization mode of GWs is maximized. For instance, a scalar longitudinal mode of GWs beyond GR has been identified as a dominant polarization component from the arm response between S/C1 and S/C4, and S/C3 and S/C4, at certain orbital positions. The GW responses between S/C4 and the arms S/C1, S/C2, and S/C3 differ from each other. As an observable quantity, $h_{\text{sum}}$ represents the amplitude of the GW response in a detector arm. $h_{\text{sum}}$ varies with the detector’s orbital position over a one-year period. The three newly added arms provide a distinct GW response compared to the traditional triangular configuration, as observed in projects such as LISA, Taiji, and Tianqin. This distinction offers additional degrees of freedom for extracting the polarization components from the total GW signals.
IV Conclusion
In conclusion, the Tetrahedron Constellation of Gravitational Wave Observatory (TEGO) is a novel system comprising four spacecrafts (S/Cs) that enhances sensitivity to gravitational wave polarizations beyond traditional ground-based or space observatories like LIGO, Virgo, KAGRA, LISA, Taiji, and TianQin. The laser telescopes and their pointing structures are mounted on the S/C platform and are evenly distributed at three locations 120 degrees apart. Comparing to a three-S/C configuration, the addition of a fourth S/C and its mounted telescope serve as redundant backups, especially for the telescopes, which together with the other two telescopes, automatically form a stable mass center for the platform without the need for trim mass. The TEGO employs more combinations of time delay interferometry configurations to effectively suppress laser frequency noise, offering more sensitivity to gravitational wave signals. As the configuration of TEGO is simultaneously sensitive to the six polarization modes of GWs, the TEGO allows for greater flexibility in extracting polarization components from total GW signals, making it simultaneously sensitive to all six polarization modes of gravitational waves.
Acknowledgements
@@
We thank the anonymous referees for very helpful suggestions, which are used to improve the manuscript. We thank Yue-Liang Wu for the theory of GW polarization beyond GR and requirement of the space mission of GW detection, which are useful to improve the manuscript. This work has been supported by the Fundamental Research Funds for the Central Universities. This work has been supported in part by the Strategic Priority Research Program of the Chinese Academy of Sciences under Grant No. XDA15020701 and XDA15020708. In this paper, the numerical computation of this work was completed at TAIJI Cluster of University of Chinese Academy of Sciences.
Author contributions
Hong-Bo Jin contributed to project designing, data preparation, analysis and manuscript writing.
Cong-Feng Qiao contributed to project designing, data preparation, analysis and manuscript writing.
All authors reviewed and commented on the final manuscript.
Competing Interests
The authors declare no competing interests.